You are evaluating two different silicon wafer milling machines. The Techron I costs $255,000, has a three-year life, and has pretax operating costs of $68,000 per year. The Techron II costs $445,000, has a five-year life, and has pretax operating costs of $41,000 per year. For both milling machines, use straight-line depreciation to zero over the project’s life and assume a salvage value of $45,000. If your tax rate is 34 percent and your discount rate is 8 percent, compute the EAC for both machines. (Your answers should be a negative value and indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EAC
Techron I $
Techron II $
Answer #1

Step 1: Calculation of NPV

-Techron I

1) Initial Cash flow (t0)

-255000

2) In between Cash flow (t1-t3)

Revenue                       -  
Less: Pretax operating cost        68,000.00
Less: Depreciation (255000/3)        85,000.00
PBT (153,000.00)
Less: Tax@34%     (52,020.00)
PAT (100,980.00)
Add: Depreciation        85,000.00
Cash flow after tax     (15,980.00)

3) Terminal Cash flow (t3)

Salvage Vale        45,000.00
Less: WDV 0
Profit on sale        45,000.00
Less: Tax @34% on above        15,300.00
Net Salvage Vale (45000-15300)        29,700.00

4) NPV

Year Cashflow PVF@8% Cashflow*PVF
0           (255,000) 1            (255,000.00)
1              (15,980) 0.9259              (14,796.30)
2              (15,980) 0.8573              (13,700.27)
3                13,720 0.7938                10,891.38

NPV = PV of Inflows - PV of Outflows

= -255000-14796.30-13700.27+10891.38

= -272605.19

-Techron II

1) Initial Cash flow (t0)

-445000

2) In between Cash flow (t1-t5)

Revenue                          -  
Less: Pretax operating cost           41,000.00
Less: Depreciation (445000/5)           89,000.00
PBT      (130,000.00)
Less: Tax@34%        (44,200.00)
PAT        (85,800.00)
Add: Depreciation           89,000.00
Cash flow after tax             3,200.00

3) Terminal Cash flow (t3)

Salvage Vale        45,000.00
Less: WDV 0
Profit on sale        45,000.00
Less: Tax @34% on above        15,300.00
Net Salvage Vale (45000-15300)        29,700.00

4) NPV

Year Cashflow PVF@8% Cashflow*PVF
0           (445,000) 1            (445,000.00)
1                  3,200 0.9259                   2,962.96
2                  3,200 0.8573                   2,743.48
3                  3,200 0.7938                   2,540.26
4                  3,200 0.7350                   2,352.10
5                32,900 0.6806                22,391.19

NPV = PV of Inflows - PV of Outflows

= (2962.96+2743.48+2540.26+2352.10+22391.19) - 445000

= 32989.99-445000

= -412010.01

Step 2: Calculation of EAC

-Techron I

EAC = NPV /(1-(1+r)^-n)/r

= -272605.19 / (1-(1+.08)^-3)/.08

= -272605.19/ 2.57709698725

= -105779.95

-Techron II

EAC = NPV /(1-(1+r)^-n)/r

= -412010.01 / (1-(1+.08)^-5)/.08

= -412010.01 / 3.99271003708

= -103190.57